OpenStudy (anonymous):

A force of 4.9 N acts on a 1 kg block kept on a frictionless surface. What is the acceleration produced by the force?

OpenStudy (irishboy123):

F = ma

OpenStudy (ooshazderoo):

you would need to do this sum \[a=\frac{ f }{ m }\] to find the acceleration sorry for any type-o's on tablet

OpenStudy (dariusx):

F = m * a a = F / m a = 4.9 / 1 = 4.9 m/s^2